Solve the following equations:
2q – 6 = 0 …...........

Grade:12

2 Answers

Harshit Singh
2 years ago
Dear Student

First we have to add 6 to the both sides of the equation,
Then, we get,
= 2q - 6 + 6 = 0 + 6
= 2q = 6
Now,
We have to divide both sides of the equation by 2,
Then, we get,
= 2q/2 = 6/2
= q = 3

Thanks
Shraddha Singh
2 years ago
Dear Student,
 
We can solve this question as follows :
2q - 6 = 0
2q = 6
q = 6/2
q = 3
 
Hope it helps!
Thank you ✌️